# HAL-Based Plugin Estimation of the Causal Dose-Response Curve

### v3

@@ 2.5 Inference: Pointwise Asymptotic Normality of Delta-Method-Based Estimators and Confidence Intervals
The HAL estimator has been shown to be asymptotically normally distributed van der Laan (2023). This property allows us to conduct inference using the delta method, providing a mechanism for constructing confidence intervals (CIs) for the estimated dose-response curve. The variance of $Q_{\beta_{n}}(x)$, where $Q_{\beta_{n}}$ is the HAL-based estimator, is estimated by the influence curve of the coefficients $\beta_{n}$.
@@
where $\Phi(a,W)$ is the basis vector for treatment $A=a$ and covariates $W$, and $IC_{\beta,P}(O_{i})=\mathbb{E}_{P}[\Phi_{i}^{T}\Phi_{i}]^{-1}\Phi_{i}^{T}(Y_{i}-\Phi_{i}^{T}\beta)$ is the influence curve of the coefficients. The empirical variance of the influence curve at each point $a$ is:
$\hat{\sigma}^{2}_{n}(a)=\frac{1}{n}\text{Var}(IC_{\Psi(a),P}(O_{i})).$
@@ 2.6 Simulation settings
Figure 1: True causal dose-response curves
@@
In each simulation, we generated 500 datasets with sample sizes of 200, 500, 1000, and 5000, following the data-generating processes outlined in Table 3 in Appendix A. These datasets were fitted by six variations of Highly Adaptive Lasso (HAL), with each model differing in terms of smoothness and regularization methods. These included: (1) zero-order smoothness with cross-validation (CV) selector, (2) zero-order smoothness with undersmoothing, (3) first-order smoothness with CV selector, (4) first-order smoothness with undersmoothing, (5) adaptive smoothness with CV selector, and (6) adaptive smoothness with undersmoothing. The zero-order models were configured using the default number of knots from the hal9001 R package Hejazi et al. (2020), catering to scenarios with indeterminate smoothness orders and knot configurations. In contrast, first-order models were trained with a smaller number of knots, suited to simulated scenarios with lower dimensionality and complexity. The adaptive models dynamically adjusted to data’s complexity, providing increased flexibility in the fitting process. For each simulation setting, oracle standard deviations were calculated across the 500 repetitions, with corresponding confidence intervals derived, to evaluate coverage under consistent estimation of the sampling variability and thus the estimators’ asymptotic normality.
These variants are evaluated based on five metrics from both perspectives of delta-method-based inferences and Monto Carlo oracle performances: coverage rates of 95% confidence intervals, bias, SE, bias/SE ratio, and mean squared error (MSE). To evaluate the selected penalties by CV and global undersmoothing criteria, comparisons were made against a grid of penalties. Furthermore, we compared the performance of globally undersmoothed smoothness adaptive HAL against established models, including the general adaptive model (GAM), polynomial regression model, and the nonparametric causal inference (npcausal) model by Kennedy et al. (2017).
@@ 3.1 CV and Undersmoothing
Table 1: Aver age pointwise performances of HAL-based plugin estimator at A = 1.0 and 4.2 on dose-response curves, n = 1000
@@
A Selector ||||Bias|||| MSE CI cov. Bias/SE SE CI cov. Bias/SE SE
@@
Figure 2: Performances in Simulation 1 and 3, n = 1000
@@ 3.2 Evaluation of Undersmoothing Criteria
Figure 3: Pointwise Performances of a grid of penalties at 5 example points in Simulation 3, n = 1000
@@ 3.3 Analysis of the Smoothness-Order Adaptive HAL
In simulation 3, interesting dynamics emerge. For smaller sample sizes (200, 500, and 1000), zero-order smoothness estimators slightly outperformed first-order estimators on the dose-response curve. However, the average CV risks were comparable or marginally higher for zero-order HAL relative to first-order HAL. As the sample size increased, first-order smoothness estimators consistently outperformed all other smoothness orders across the evaluation metrics. Furthermore, the smoothness level selected by the discrete Super Learner adapted to the sample size of the training data. For smaller sample sizes of 200, 500, and 1000, the HAL chose first-order smoothness 56%, 62.1%, and 67.4% of the time, respectively. For the largest sample size of 5000, the HAL demonstrated a definitive preference, selecting first-order smoothness in all iterations (100%). Such adaptive choices of smoothness order consistently align with the smoothness order that performed best for the available data.
@@ 3.4 Comparison with Other Methods
Figure 4: Oracle Performances of 4 estimators in 4 simulations, n = 5000
Table 2: Average Performances of 4 estimators in 4 simulations, n = 5000
@@
Method ||||Bias|||| MSE CI cov. Bias/SE SE CI cov. Bias/SE SE
@@ 4 Discussions
In this paper, we explored the performance of different methods for dose-response estimation through a series of simulations, with particular emphasis on the Highly Adaptive Lasso (HAL) plug-in estimator. We investigated the reliability of the HAL estimator and examined critical estimation criteria, including penalty selection and smoothness order choices.
@@
Our results highlight the importance of choice between cross-validation (CV) and undersmoothing criteria when selecting the penalty for HAL. In simple sinarios, such as in simulation 1, where the dose-response curve is smooth, CV and undersmoothing approaches exhibit comparable performances. However, as curve complexity increases, such as in simulation 3, the undersmoothing approach stands out by achieving higher oracle coverage rates and lower biases. This finding indicates that undersmoothing is particularly beneficial for accurately capturing complex dose-response relationships, striking an optimal balance between bias and variance. Importantly, undersmoothing enables HAL to adjust across various treatment levels on the curve effectively. Additionally, the adaptive smoothness selects the optimal smoothness orders given the dose-response curve’s complexity and available sample size. With big enough sample sizes, first-order HAL demonstrates the flexibility to handle both simple and complex (non-smooth) curve structures.
@@
When comparing the four estimators—HAL, Generalized Additive Models (GAM), polynomial regression, and the nonparametric causal inference approach (npcausal)—our results show that the HAL-based plug-in estimator consistently delivers superior performance. It achieves the highest CI coverage rates, lowest absolute biases, and smallest mean squared error (MSE) values, especially in scenarios with more complex dose-response curves. In contrast, other estimators, particularly the polynomial model, show considerable variability and fail to maintain the same level of performance. These results strongly suggest that the HAL-based plug-in estimator is a preferred method for dose-response estimation due to its robustness, accuracy, and flexibility in capturing the true underlying patterns within the data without leaning on strong model specification assumptions.
5 Conclusions and Future Work
@@
In this study, we evaluated the effectiveness of the HAL-based plug-in estimator for estimating causal dose-response curves across various simulation scenarios. Our findings demonstrate that the HAL-based estimator consistently outperforms traditional methods, such as GAM and polynomial regression, as well as more advanced approaches like the npcausal method introduced by Kennedy et al. (2017). The estimator’s ability to provide unbiased estimates of complex dose-response relationships and deliver confidence intervals with nominal coverage (valid inference) further enhances its robustness. Its adaptability to increasing complexity and varying sample sizes, facilitated by smoothness-order adjustments, highlights its potential for broad application in diverse practical contexts.
@@
While HAL shows considerable promise in modeling dose-response curves, there are computational challenges. Its implementation requires substantial computing time and memory, as a compromise of including a large set of basis functions that flexibly fit a wide range of true curves without assuming specific parametric forms. Solving the inverse of a matrix with many basis functions, especially those selected with the undersmoothing criteria, may lead to invertibility issues. Ongoing research aims to address these concerns by improving HAL’s scalability and computational efficiency.
@@
Our approach shows promise for making inferences on other nonpathwise differentiable parameters, such as the conditional average treatment effect (CATE) by HAL. Moreover, advances in the adaptive targeted maximum likelihood framework van der Laan et al. (2024) may further improve robustness for nonpathwise differentiable functions. These developments, coupled with the adoption of high-dimensional parametric working models revealed by an initial HAL fit, suggest substantial future improvements in estimation and inference for complex causal parameters.
The source code for our methods is available at https://github.com/SeraphinaShi/HAL-DoseResponseCurve, supporting reproducibility and encouraging further research in this area.
Appendix A
A.1 A. Data Generating Distributions (DGDs) in 4 Simulations
@@
A.2 B. Compare performances of estimators based on HAL with different smoothness order splines

### v4

@@ 1.5 Inference: Pointwise Asymptotic Normality of Delta-Method-Based Estimators and Confidence Intervals
The HAL estimator has been shown to be asymptotically normally distributed (van der Laan, 2023). This property allows us to conduct inference using the delta method, providing a mechanism for constructing confidence intervals (CIs) for the estimated dose-response curve. The variance of $Q_{\beta_{n}}(x)$, where $Q_{\beta_{n}}$ is the HAL-based estimator, is estimated by the influence curve of the coefficients $\beta_{n}$.
@@
where $\Phi(a,W)$ is the basis vector for treatment $A=a$ and covariates $W$, and $IC_{\beta,P}(O_{i})=\mathbb{E}_{P}[\Phi_{i}^{T}\Phi_{i}]^{-1}\Phi_{i}^{T}(Y_{i}-\Phi_{i}^{T}\beta)$ is the influence curve of the coefficients. The variance of the estimator $\hat{\Psi}_{n}(a)$ is estimated by computing the empirical variance of the influence curve:
$\hat{\sigma}^{2}_{n}(a)=\frac{1}{n}\hat{\text{Var}}(IC_{\Psi(a),P}(O_{i})),$
where $\hat{\text{Var}}$ denotes the sample variance computed from the estimated influence curve evaluated on the observed data. Because $\hat{\sigma}_{n}^{2}(a)$ estimates the variance of $\hat{\Psi}_{n}(a)$, it can be directly used to construct confidence intervals.
@@ 1.6 Simulation settings
Figure 1: True causal dose-response curves across four simulations. The ground-truth dose-response relationships exhibit increasing complexity, ranging from monotonic trends (Simulation 1) to highly non-linear and discontinuous patterns (Simulations 2–4).
@@
In each simulation, we generated 500 datasets with sample sizes of 200, 500, 1000, and 5000, following the data-generating processes outlined in Table  3 in Appendix A. These datasets were fitted by six variations of Highly Adaptive Lasso (HAL), with each model differing in terms of smoothness and regularization methods. These included: (1) zero-order smoothness with cross-validation (CV) selector, (2) zero-order smoothness with undersmoothing, (3) first-order smoothness with CV selector, (4) first-order smoothness with undersmoothing, (5) adaptive smoothness with CV selector, and (6) adaptive smoothness with undersmoothing. The zero-order models were configured using the default number of knots from the hal9001 R package (Hejazi et al., 2020), catering to scenarios with indeterminate smoothness orders and knot configurations. In contrast, first-order models were trained with a smaller number of knots, suited to simulated scenarios with lower dimensionality and complexity. The adaptive models dynamically adjusted to the data’s complexity, providing increased flexibility in the fitting process. For each simulation setting, oracle standard deviations were calculated across the 500 repetitions, with corresponding confidence intervals derived, to evaluate coverage under consistent estimation of the sampling variability and thus the estimators’ asymptotic normality.
These variants are evaluated based on five metrics from both perspectives of delta-method-based inferences and Monte Carlo oracle performances: coverage rates of 95% confidence intervals, bias, SE, bias/SE ratio, and mean squared error (MSE). To estimate coverage, we report both that based on the estimated delta-method-based standard error and on the oracle standard error, calculated as the standard deviation of the estimator over the repeated simulations. We did this to separate biases in estimating the standard error from non-normal finite sampling distributions. To evaluate the selected penalties by CV and global undersmoothing criteria, comparisons were made against a grid of penalties. Furthermore, we compared the performance of globally undersmoothed smoothness adaptive HAL against established models, including the general adaptive model (GAM), polynomial regression model, and the nonparametric causal inference (npcausal) model by Kennedy et al. (2017).
@@ 2.1 CV and Undersmoothing
Table 1: Average pointwise performance of the HAL-based plug-in estimator at $A=1.0$ and $A=4.2$ across four simulation settings (n = 1000). The table compares the performance of the HAL estimator under two selection strategies: cross-validation (CV) and undersmoothing (U). For each simulation scenario and evaluation point, we report bias, mean squared error (MSE), confidence interval (CI) coverage, bias-to-standard error ratio (Bias/SE), and standard error (SE), using both delta-method-based and oracle CI constructions. In general, undersmoothing improves CI coverage and yields smaller Bias/SE ratios, particularly in more complex simulation settings (Simulations 3 and 4). However, this comes with a slight trade-off in increased standard error. The oracle results serve as a benchmark for evaluating the accuracy of the delta-method-based inference.
@@
A Selector ||Bias|| MSE CI cov. Bias/SE SE CI cov. Bias/SE SE
@@
Figure 2: Performance comparison in Simulation 1 and Simulation 3 (n = 1000). The blue lines represent estimated dose-response curves and corresponding 95% confidence intervals (CIs) using cross-validated HAL (CV-HAL), while the orange lines represent those using undersmoothed HAL. Solid lines denote the oracle CIs derived from 500 Monte Carlo repetitions, and dashed lines indicate delta-method-based CIs. For the simpler curve in Simulation 1, both CV-HAL and undersmoothed HAL perform similarly. In contrast, for the more complex curve in Simulation 3, undersmoothing improves performance by yielding higher CI coverage rates and lower mean squared error (MSE). Additionally, delta-method-based CIs closely approximate the oracle CIs, demonstrating their accuracy.
@@ 2.2 Evaluation of Undersmoothing Criteria
Figure 3: Pointwise performance across a grid of penalty values at five example locations in Simulation 3 (n = 1000). Each column corresponds to a target point along the dose-response curve, with increasing complexity from left to right. The rows display: (1) estimated values and 95% CIs; (2) CI coverage rates; (3) log bias-to-standard error ratio; (4) absolute bias; (5) standard error; and (6) mean squared error (MSE). Results aggregated over 500 Monte Carlo repetitions are shown in green (oracle-based) and pink (delta-method-based). The dashed blue vertical lines indicate the penalty selected by cross-validation (CV), and the dashed orange lines indicate the penalty selected by the undersmoothing criterion. The undersmoothing approach consistently selects smaller penalties and often strikes a better balance across bias, variance, and MSE, particularly in regions with higher curvature.
@@ 2.3 Analysis of the Smoothness-Order Adaptive HAL
In simulation 3, interesting dynamics emerge. For smaller sample sizes (200, 500, and 1000), zero-order HAL slightly outperformed first-order HAL on the dose-response curve. However, the average CV risks were comparable or marginally higher for zero-order HAL relative to first-order HAL. As the sample size increased, first-order smoothness estimators consistently outperformed all other smoothness orders across the evaluation metrics. Furthermore, the smoothness level selected by the discrete Super Learner adapted to the sample size of the training data. For smaller sample sizes of 200, 500, and 1000, the method based upon adaptive selection of spline order chose first-order smoothness 56%, 62.1%, and 67.4% of the time, respectively. For the largest sample size of 5000, the HAL demonstrated a definitive preference, selecting first-order smoothness in all iterations (100%). Such adaptive choices of smoothness order consistently align with the smoothness order that performed best for the available data.
@@ 2.4 Comparison with Other Methods
Figure 4: Oracle performance of four estimators across four simulation settings (n = 5000). Each column corresponds to a simulation scenario with increasing complexity from left to right. The rows display: (1) point estimates and 95% confidence intervals (CIs), (2) CI coverage rates, (3) log bias-to-standard error (SE) ratio, and (4) mean squared error (MSE). The four estimators include HAL (green), GAM (orange), polynomial regression (pink), and nonparametric causal regression (npcausal, purple). In Simulation 1, all methods approximate the true curve well, with HAL and npcausal slightly outperforming the others. In Simulations 2–4, polynomial regression fails to capture the complex curve shapes, while HAL consistently demonstrates superior performance across all metrics, including accuracy, CI coverage, and MSE.
Table 2: Average performance of four estimators across four simulation settings (n = 5000). This table summarizes the average bias, mean squared error (MSE), confidence interval (CI) coverage, bias-to-standard error ratio (Bias/SE), and standard error (SE) for four methods: Highly Adaptive Lasso (HAL), nonparametric causal regression (npcausal), generalized additive models (GAM), and polynomial regression (Poly). Results are reported for both oracle-based and delta-method-based inference procedures. HAL consistently achieves the best overall performance, with low bias, high CI coverage, and balanced Bias/SE across all simulation scenarios. While npcausal performs competitively in some settings, both GAM and polynomial regression exhibit deteriorating performance as the complexity of the data-generating process increases, particularly in Simulations 3 and 4. Polynomial regression fails to maintain accurate CI coverage and has the highest Bias/SE ratios and MSEs under complex scenarios.
@@
Method ||Bias|| MSE CI cov. Bias/SE SE CI cov. Bias/SE SE
@@ 3 Discussions
In this paper, we explored the performance of different methods for causal dose-response estimation through a series of simulations, with a novel emphasis on the Highly Adaptive Lasso (HAL) plug-in estimator. Our contribution lies in implementing HAL with inference in the context of dose-response estimation and systematically evaluating its empirical robustness and adaptability across a range of complex scenarios, areas which have not been thoroughly addressed in previous work. We investigated the reliability of the HAL estimator and examined estimation criteria, including penalty selection and smoothness order choices.
@@
Our results highlight the importance the of choice between cross-validation (CV) and undersmoothing when selecting HAL’s penalty parameter. In simple scenarios, such as in simulation 1, where the dose-response curve is smooth, CV and undersmoothing approaches exhibit comparable performances. However, as curve complexity increases, such as in simulation 3, the undersmoothing approach stands out by achieving higher oracle coverage rates and lower biases. This finding indicates that undersmoothing is particularly beneficial for accurately capturing complex dose-response relationships, striking an optimal balance between bias and variance. Importantly, undersmoothing enables HAL to adjust across various treatment levels on the curve effectively. Additionally, smoothness-adaptive HAL selects the optimal smoothness orders given the dose-response curve’s complexity and available sample size. With sufficiently large sample sizes, first-order HAL demonstrates the flexibility to handle both simple and complex (non-smooth) curve structures.
@@
When comparing the four estimators—HAL, Generalized Additive Models (GAM), polynomial regression, and the nonparametric causal inference approach (npcausal)—our results show that the HAL-based plug-in estimator consistently delivers superior performance. It achieves the highest CI coverage rates, lowest absolute biases, and smallest mean squared error (MSE) values, especially in scenarios with more complex dose-response curves. In contrast, other estimators, particularly the polynomial model, show considerable variability and fail to maintain the same level of performance. These findings establish the HAL-based plug-in estimator as a preferred method for dose-response estimation, due to its robustness, accuracy, and ability to capture underlying relationships without relying on restrictive parametric assumptions.
@@
The estimator’s ability to provide unbiased estimates of complex dose-response relationships and deliver confidence intervals achieving nominal coverage (valid inference) further enhances its robustness and practical applicability. Its adaptability to increasing complexity and varying sample sizes, facilitated by smoothness-order adjustments, highlights its potential for broad application in diverse practical contexts.
@@
While HAL shows considerable promise in modeling dose-response curves, there are computational challenges. Specifically, its implementation demands substantial computing resources, both in terms of time and memory, as it incorporates a comprehensive set of basis functions to flexibly fit a wide range of true curves without assuming specific parametric forms. Moreover, solving the inverse of a matrix with many basis functions, especially those selected with the undersmoothing criteria, may lead to invertibility issues. Ongoing research aims to address these concerns by improving HAL’s scalability and computational efficiency.
@@
Our approach presents HAL’s capability of providing statistical inference for non-pathwise differentiable parameters, as demonstrated for the conditional mean and the conditional average treatment effect (Zhang et al., 2025; Zhang et al., 2024). A brief overview of this capability is also provided in Butzin-Dozier et al. (2024). Moreover, advances in the adaptive targeted maximum likelihood framework (van der Laan et al., 2024) may further improve robustness for nonpathwise differentiable functions. These developments, coupled with the adoption of high-dimensional parametric working models revealed by an initial HAL fit, suggest substantial future improvements in estimation and inference for complex causal parameters.
4 Conclusion
In this paper, we presented the HAL-based plug-in estimator for continuous causal dose-response curves, highlighting both its theoretical underpinnings and empirical performance. By leveraging the flexibility of càdlàg functions and the adaptivity of HAL, the estimator achieves pointwise asymptotic normality under weak conditions, even for parameters that are non-pathwise differentiable.. Through extensive simulations, we demonstrated its robustness across diverse settings, consistently yielding valid coverage, reduced bias, and superior accuracy relative to established alternatives such as GAM, polynomial regression, and npcausal estimators. Importantly, undersmoothing and smoothness-adaptive selection further enhance performance, enabling the method to adapt to increasing complexity with growing sample sizes.
Together, these findings position the HAL plug-in estimator as a practical, theoretically grounded, and empirically validated approach for dose-response estimation in continuous treatment settings, bridging the gap between asymptotic theory and applied causal inference.
The halDRC R package, which includes both real-data application examples and simulated sample demos, is available at https://github.com/SeraphinaShi/halDRC. The source code implementing our methods can be found at https://github.com/SeraphinaShi/HAL-DoseResponseCurve. These resources support reproducibility and are intended to facilitate further research in this area.
Appendix
A. Data Generating Distributions (DGDs) in 4 Simulations
@@
B. Compare performances of estimators based on HAL with different smoothness order splines